# Jamais Vu: Exposing the Generalization Gap in Supervised Semantic Correspondence

### v1

@@ 3.3 Dense Geometric Alignment
This is equivalent to computing descriptors via softmax attention over $\mathcal{Z}$, using $\Phi(I,\bm{u})$ as queries, $\bm{z}$ as keys, and $\bm{p}$ as values. For labeled keypoints $\bm{k}_{l}$, minimizing Eq. 4 ensures $\varphi(I,\bm{k}_{l})=\bm{p}_{l}$.
@@
For each training image $I$, we can now estimate the dense canonical coordinates $\mathcal{X}^{c}$ over its pixels via Eq. 5, and the posed coordinates $\mathcal{X}^{b}$ via depth backprojection using Eq. 2. In practice, $\mathcal{X}^{c}$ and $\mathcal{X}^{b}$ only consist of object points that are selected using the segmentation mask. We aim to align these two representations so that $\mathcal{X}^{c}$ properly reflects the object geometry. However, our annotations are only sparse, thus we cannot directly supervise $\varphi(I,\bm{u})$ for arbitrary coordinates $\bm{u}$. Instead, we make the assumption that even though the posed and canonical shape are different, they should be locally similar. We encourage the geometric alignment between a small neighborhood of points sampled in the posed space, and their corresponding locations in the canonical space.
@@
For a given point $q\in\mathcal{X}^{c}$, we sample its $k$ nearest neighbors to obtain $\mathbf{C}_{q}$ in the canonical space, and the corresponding coordinates in the posed space $\mathbf{B}_{q}$, and minimize the alignment error between two sets. We also sample neighbors of a given point $r\in\mathcal{X}^{b}$ in the posed space, denoted as $\mathbf{B}_{r}$, and compute the loss in the other direction:
$\min_{\Phi,\mathcal{Z}}\frac{1}{N}\sum_{n=1}^{N}||\mathbf{C}_{q}^{(n)}-\mathbf{M}_{c2b}^{(n)}\mathbf{B}_{q}^{(n)}||_{1}+||\mathbf{B}_{r}^{(n)}-\mathbf{M}_{b2c}^{(n)}\mathbf{C}_{r}^{(n)}||_{1},$ (6)
@@
where $\mathbf{M}_{c2b}$ and $\mathbf{M}_{b2c}$ are the rigid transformations between the canonical and posed coordinates, computed using Kabsch’s algorithm at each iteration as in Eq. 3.
@@
In the canonical space, neighbors are selected using a standard k-nearest neighbors strategy. However, this approach can be unreliable in the posed space due to object deformations. For instance, in the case of a person eating, the hand might be close to the face in 3D space, and thus points belonging to the face might mistakenly be selected as neighbors of the hand. Instead, we use a geodesic sampling strategy, that samples points along the surface of the object. Starting from a seed point, we iteratively grow the neighborhood by selecting the next point with the shortest geodesic distance to the current set, effectively approximating surface-based proximity rather than raw spatial closeness.
@@
We jointly optimize the descriptor learning loss in Eq. 4 and the geometric consistency loss in Eq. 6 to learn $\Phi$ and $\mathcal{P}$. While these objectives suffice to learn an SC model, in practice we build our implementation on Geo-SC zhang2023telling and optimize its parameters jointly over the sum of our objective and the original one. At inference time, rather than simply querying nearest-neighbor in the descriptor space predicted by $\Phi$, we make use of the soft-argmax window matching strategy proposed in zhang2023telling . Unlike $\varphi$, which relies on category-specific canonical coordinate set $\mathcal{P}$ and descriptors $\mathcal{Z}$, $\Phi$ can be applied to previously unseen object categories directly.
@@ 4 SPair-U: A Benchmark for Evaluating Unseen Keypoints
Figure 3: Example keypoint annotations contained within our new SPair-U evaluation dataset. It utilizes the same images as the SPair-71k dataset min2019spair , but adds additional keypoints not present in SPair-71k. This enables benchmarking of SC methods on the existing keypoints along with our new ones. On the right we summarize the main statistics of our new dataset.
As illustrated in Fig. 1, the performance of the state-of-the-art supervised SC methods zhang2023telling ; taleof2feats degrades significantly when queried on keypoints that are not part of the training set. We posit that this is caused by models only learning strong representations for these specific points, while largely ignoring the remaining pixels. We would like to assess the performance of SC methods when evaluated on keypoints that were previously unseen (i.e., not in the labeled set) at training time. A possible solution is to use an existing dataset while splitting the annotations into two mutually exclusive sets of keypoints, seen and unseen, between training and evaluation. However, this strategy would reduce the supervision available, and require retraining previous techniques for evaluation.
Instead, we introduce a new evaluation benchmark, SPair-U, by labeling additional keypoints from the SPair-71k dataset min2019spair . We added at least four new points for each of the 18 categories found in SPair-71k. For animals, we focused on additional joints on the limbs, and for vehicles we added semantic parts that were not already labeled, e.g., windshield or fenders. Boats, bottles, potted plants, and tv monitors keypoints are not semantic per se in SPair-71k, but are rather spread around on the outline of the object. Thus, we added midway points between those already defined. In total, we add 1,272 new test keypoint annotations resulting in 8,254 new test pairs. We illustrate some of these new annotations in Fig. 3, and the full list with more details can be found in Appendix D. As shown in Fig. 1, current supervised methods tend to default to making predictions new training keypoint locations when predicting correspondences for the SPair-U keypoints.
@@ 5.1 Implementation Details
Our 3D prototype design being independent from typical semantic correspondence architectures, we can add it as an additional objective on top of established models. Thus, we base our experiments on Geo-SC zhang2023telling , strictly following provided hyperparameters, e.g., learning rate, batch size, optimizer, scheduler, epoch count, simply adding our additional loss terms and jointly optimizing $\mathcal{P}$ and $\mathcal{Z}$ alongside Geo-SC’s feature extractor $\Phi$. In practice, $\Phi$ consists of additional bottleneck layers trained on top of frozen DINOv2 and SD backbones. We extract depth maps and camera intrinsics using MoGe wang2024moge , and use Segment Anything kirillov2023segment to obtain segmentation masks. Additional implementation details are provided in Appendix A.
@@ 5.2 Quantitative Results
SPair-71k. We first compare our models to other SC approaches on the SPair-71k benchmark min2019spair , which contains images from 18 categories. We use the standard PCK@$0.1_{bbox}$ which considers a match to be correct if its prediction lies within distance $0.1\times\max(h,w)$ of the ground truth location where $(h,w)$ is the height and width of the target object bounding box. Typically, supervised models report per-image PCK, i.e., the average score of each image per category, while unsupervised ones use per point PCK, i.e., the average number of correct matches per category. In order to properly compare performances between the two families, we recompute per-image PCK for all models, which results in a small drop for the unsupervised models.
Results in Table 1 show that our model ranks competitively against other approaches, with a marginal $0.2\%$ performance drop on average against its backbone Geo-SC zhang2023telling . Per category results show small improvements in 9 of the categories, the highest one being $3.6\%$ on bus, and small drop on the other 9, the largest being $2.4\%$ on bottle. Overall, the differences are minor, illustrating that adding our extra objective does not interfere with the original model.
Table 1: Results on standard evaluation keypoints for SPair-71k. Per-image PCK@$0.1_{bbox}$ scores are reported. In this table and the following: All models use the soft matching strategy described in zhang2023telling except those followed by *. Models with a dagger benefit from AP-10K pretraining. Models in the $\mathcal{K}$ category use keypoint supervision, while $\mathcal{K}$ do not. Best results are bolded and second best are underlined.
@@
Table 2: Results on unseen keypoints on our SPair-U benchmark. Per-image PCK@$0.1_{bbox}$ scores on unseen keypoints are reported.
@@
SPair-U. To evaluate the model’s ability to generalize to unseen semantic points, we assess its performance on the introduced SPair-U keypoints using per-image PCK@0.1bbox for a like-for-like comparison. We exclude the Test-time Adaptive Pose Alignment from zhang2023telling since it requires prior knowledge of keypoint semantics to relabel flipped keypoints, which contradicts the assumption that evaluation keypoints are unknown.
As shown in Table 2, results on SPair-U reveal a stark contrast between supervised and unsupervised models. While unsupervised models see only a modest performance drop, likely due to increased task difficulty, supervised models experience a significant decline. Many of the pre-existing approaches are outperformed by the unsupervised DINO+SD baseline taleof2feats and they are consistently beaten by the weakly-supervised Spherical Maps mariotti2024improving . Notably, in 8 categories, the best-performing model is unsupervised, suggesting that supervised approaches behave more like keypoint regressors and fail to generalize to novel correspondences.
@@
As shown in Table 3, while Geo-SC zhang2023telling achieves the best performance on the standard SPair-71k test set, it underperforms on all other benchmarks, highlighting its limited generalization. Notably, even a simple supervised DINO+SD taleof2feats baseline outperforms Geo-SC at the standard $0.10$ threshold when using the same soft window matching strategy. This stands in sharp contrast to the findings in zhang2023telling where Geo-SC consistently outperforms its baseline by $10\%$ on the 3 AP-10K benchmarks, when both models are trained on AP-10K, indicating potential overfitting to that dataset.
@@ 5.3 Qualitative Results
Figure 4: PCA decomposition of the feature maps for different models. Note that PCA is computed on object features only. The inclusion of geometric constraints during training results in fewer high frequency artifacts in the predicted feature maps for our approach.
In Fig. 4, we visualize Principal Component Analysis of object features produced by our model, Geo-SC and the unsupervised DINO+SD. Our model produces descriptors that vary smoothly over the object surface while uniquely identifying each point. In comparison, the predictions of Geo-SC are noisier, with sudden discontinuities (cf. bus) and uniform descriptors on regions that have no keypoints (cf. the body of the cat and the cow). Meanwhile, the unsupervised features fail to separate repeated parts (cf. plane engines) and produces noisy features in textureless areas (cf. tv).
Following  wandel2025semalign3d , we visualize our learned canonical shapes in Fig. 5 by randomly logging predicted canonical coordinates of objects in order to overlap multiple partial point clouds over the training data. We observe that the spatial organization of $\mathcal{P}$, i.e., the large bold points, captures the general shape of the category, and the predicted coordinates densely span the object surface. Interestingly, our parametrization of the canonical shape Eq. 5 forces predicted coordinates to lie within the convex hull of $\mathcal{P}$, which explains the incomplete wheel on the motorbike. We also observe that very few points are mapped towards the end of the train, which we attribute to the varying length of trains across instances and the bias toward frontal viewpoints. Note that contrary to wandel2025semalign3d , these are simply visualizations and are not used for inference, meaning this limitation is unlikely to significantly affect performance.
@@ 6 Limitations
While augmenting the test set of SPair-71k with new keypoints enables us to evaluate existing techniques with their provided models and code, our proposed benchmark SPair-U inherits some drawbacks from SPair-71k. In particular, the test set is small, consisting of only 481 images but with over 8,000 test pairs, and the categories are restricted to only common objects and animals. Furthermore, few categories were already labeled with a high number of keypoints where it is potentially easier to detect the newly added ones by relating them to the existing ones. While our findings remain valid about the generalization issues in supervised SC techniques, a larger, higher-quality held-out set of images and points would be beneficial for more extensive evaluation.
@@
Finally, our assumption that geometry is a good proxy for semantics breaks down for complex object categories with diverse spatial part configurations. For example, cabinets may have different numbers of doors that open in various directions, leading to inconsistent placement of features like handles. Additionally, we do not foresee any negative social impacts of our work.
@@ 7 Conclusion
We tackled the problem of estimating semantic correspondences between different instances of the same object category. While recent supervised methods have made significant progress on known keypoints, we showed that their representations often fail to generalize to unseen keypoints not annotated during training. To address this limitation, we proposed a new approach that incorporates geometric constraints at training time through a per-category, learned continuous canonical manifold. Our method achieves superior performance compared to both supervised and unsupervised baselines, as demonstrated on SPair-U, a new dataset we introduce with additional keypoint annotations for the widely used SPair-71k benchmark.
Acknowledgments. This project was supported by the EPSRC Visual AI grant EP/T028572/1.

### v2

@@ 3.3 Dense Geometric Alignment
This is equivalent to computing descriptors via softmax attention over $\mathcal{Z}$, using $\Phi(I,\bm{u})$ as queries, $\bm{z}$ as keys, and $\bm{p}$ as values. For labeled keypoints $\bm{k}_{l}$, minimizing Eq.˜4 ensures $\varphi(I,\bm{k}_{l})=\bm{p}_{l}$.
@@
For each training image $I$, we can now estimate the dense canonical coordinates $\mathcal{X}^{c}$ over its pixels via Eq.˜5, and the posed coordinates $\mathcal{X}^{b}$ via depth backprojection using Eq.˜2. In practice, $\mathcal{X}^{c}$ and $\mathcal{X}^{b}$ only consist of object points that are selected using an object segmentation mask. We aim to align these two representations so that $\mathcal{X}^{c}$ properly reflects the object geometry. However, our annotations are only sparse, thus we cannot directly supervise $\varphi(I,\bm{u})$ for arbitrary coordinates $\bm{u}$. Instead, we make the assumption that even though the posed and canonical shape are different, they should be locally similar. We encourage the geometric alignment between a small neighborhood of points sampled in the posed space, and their corresponding locations in the canonical space.
@@
For a given point $\mathbf{q}\in\mathcal{X}^{c}$, we sample its $k$ nearest neighbors to obtain $\mathbf{C}_{q}$ in the canonical space, and the corresponding coordinates in the posed space $\mathbf{B}_{q}$, and minimize the alignment error between two sets. We also sample neighbors of a given point $r\in\mathcal{X}^{b}$ in the posed space, denoted as $\mathbf{B}_{r}$, and compute the loss in the other direction:
[…]
@@
where $\mathbf{M}_{c2b}$ and $\mathbf{M}_{b2c}$ are the rigid transformations between the canonical and posed coordinates, computed using the Kabsch-Umeyama algorithm at each iteration as in Eq.˜3.
@@
In the canonical space, neighbors are selected using a standard k-nearest neighbors strategy. However, this approach can be unreliable in the posed space due to object deformations. For instance, in the case of a person eating, the hand might be close to the face in 3D space, and thus points belonging to the face might mistakenly be selected as neighbors of the hand. Instead, we use a pseudo-geodesic sampling strategy that samples points along the surface of the object. Starting from a seed point, we iteratively grow the neighborhood by selecting the next point with the shortest distance to the current set, effectively approximating surface-based proximity rather than raw spatial closeness. Pseudocode is provided in Algorithm˜1.
@@
We jointly optimize the descriptor learning loss in Eq.˜4 and the geometric consistency loss in Eq.˜6 to learn $\Phi$ and $\mathcal{P}$. While these objectives suffice to learn a SC model, in practice we build our implementation on Geo-SC zhang2023telling and optimize its parameters jointly over the sum of our objective and the original one. At inference time, rather than simply querying nearest-neighbor in the descriptor space predicted by $\Phi$, we make use of the soft-argmax window matching strategy proposed in zhang2023telling. Unlike $\varphi$, which relies on category-specific canonical coordinate set $\mathcal{P}$ and descriptors $\mathcal{Z}$, $\Phi$ can be applied to previously unseen object categories directly.
@@ 4 SPair-U: A Benchmark for Evaluating Unseen Keypoints
Figure 3: Example keypoint annotations from our new SPair-U evaluation dataset. It utilizes the same images as the SPair-71k dataset min2019spair, but adds additional keypoints not present in SPair-71k. This enables benchmarking of SC methods on the existing keypoints along with our new ones. On the right we summarize the main statistics of our new dataset.
As illustrated in Fig.˜1, the performance of the state-of-the-art supervised SC methods zhang2023telling; taleof2feats degrades significantly when queried on keypoints that are not part of their training sets. We posit that this is caused by models only learning strong representations for these specific points, while largely ignoring the remaining pixels. We would like to assess the performance of SC methods when evaluated on keypoints that were previously unseen (i.e., not in the labeled set) at training time. A possible solution is to use an existing dataset while splitting the annotations into two mutually exclusive sets of keypoints, seen and unseen, between training and evaluation. However, this strategy would reduce the supervision available, and require retraining previous techniques for evaluation.
Instead, we introduce a new evaluation benchmark, SPair-U, by labeling additional keypoints from the SPair-71k dataset min2019spair. We added at least four new points for each of the 18 categories found in SPair-71k. For animals, we focused on additional joints on the limbs, and for vehicles we added semantic parts that were not already labeled, e.g., windshield or fenders. Boats, bottles, potted plants, and tv monitors keypoints are not semantic per se in SPair-71k, but are rather spread around on the outline of the object. Thus, we added midway points between those already defined. In total, we add 1,272 new individual test keypoint annotations resulting in 19,990 new keypoint pairs spread across 8,254 image pairs. We illustrate some of these new annotations in Fig.˜3, and the full list with more details can be found in LABEL:{tab:kp_desc}. As shown in Fig.˜1, current supervised methods tend to predict locations of keypoints seen during when queried on the new SPair-U points.
@@ 5.1 Implementation Details
Our 3D prototype approach is complementary to existing semantic correspondence architectures, thus we can add it as an additional objective on top of established models. We base our experiments on Geo-SC zhang2023telling, strictly following their provided hyperparameters, e.g., learning rate, batch size, optimizer, scheduler, and epoch count, simply adding our additional loss terms and jointly optimizing $\mathcal{P}$ and $\mathcal{Z}$ alongside Geo-SC’s feature extractor $\Phi$. We also preserve the contrastive $\mathcal{L}_{\text{sparse}}$ and dense $\mathcal{L}_{\text{dense}}$ objectives with gaussian noise, as well as feature maps dropout and pose-variant augmentation. We refer to the original publication and official implementation for in-depth description of these features. While $\mathcal{P}$ and $\mathcal{Z}$ are category-specific, a single $\Phi$ is trained on the full dataset, allowing generalization to new categories. Furthermore, gradients coming from Eq.˜6 are not backpropagated to $\mathcal{P}$, meaning it is only optimized using Eq.˜3.
Our complete loss term is $\mathcal{L}_{\text{sparse}}+\mathcal{L}_{\text{dense}}+\mathcal{L}_{\mathcal{P}}+0.3\times\mathcal{L}_{\mathcal{Z}}+\mathcal{L}_{\text{geom}}$, where $\mathcal{L}_{\mathcal{P}}$, $\mathcal{L}_{\mathcal{Z}}$, and $\mathcal{L}_{\text{geom}}$ correspond to Eq.˜3, Eq.˜4 and Eq.˜6 respectively. The justification for setting the weight $\lambda_{\mathcal{Z}}=0.3$ is provided in Appendix˜B.
In practice, $\Phi$ consists of additional bottleneck layers trained on top of frozen DINOv2 and SD backbones. We extract depth maps and camera intrinsics using MoGe wang2024moge, and use Segment Anything kirillov2023segment to obtain segmentation masks. Importantly, these are only used during training. During evaluation, matches are computed only from the predictions of $\Phi$. Additional implementation details can be found in Appendix˜A.
@@ 5.2 Quantitative Results
M v [Uncaptioned image] Ω T [Uncaptioned image] [Uncaptioned image] avg
@@
M v [Uncaptioned image] Ω T [Uncaptioned image] [Uncaptioned image] avg
@@
[…]
Results on seen keypoints in Table˜1 show that our model ranks competitively against other approaches, with a marginal $0.2\%$ performance drop on average against its backbone Geo-SC zhang2023telling. Per-category results show small improvements in nine of the categories, the highest one being $3.6\%$ on bus, and small drop on the other nine, the largest being $2.4\%$ on bottle. Overall, the differences are minor, illustrating that adding our extra objective does not interfere with the original model.
Unseen – SPair-U. To evaluate a model’s ability to generalize to unseen semantic points, we assess its performance on our new SPair-U keypoints using per-image PCK@0.1bbox for a like-for-like comparison. We exclude the Test-time Adaptive Pose Alignment from zhang2023telling since it requires prior knowledge of keypoint semantics to relabel flipped keypoints, which contradicts the assumption that evaluation keypoints are unknown.
As shown in Table˜2, results on SPair-U reveal a stark contrast between supervised and unsupervised models. While unsupervised models see only a modest performance drop, likely due to increased task difficulty, supervised models experience a significant decline. Many of the pre-existing approaches are outperformed by the unsupervised DINO+SD baseline taleof2feats and they are consistently beaten by the weakly-supervised Spherical Maps mariotti2024improving. Notably, in eight categories, the best-performing model Has not seen keypoints during training, suggesting that supervised approaches behave more like keypoint regressors and fail to generalize to novel correspondences.
@@
As shown in Table˜3, while Geo-SC zhang2023telling achieves the best performance on the standard SPair-71k test set, it underperforms on all other benchmarks, highlighting its limited generalization. Notably, even a simple supervised DINO+SD taleof2feats baseline outperforms Geo-SC at the standard $0.10$ threshold when using the same soft window matching strategy. This stands in sharp contrast to the findings in zhang2023telling where Geo-SC consistently outperforms its baseline by $10\%$ on the three AP-10K benchmarks, when both models are trained on AP-10K, indicating potential overfitting to that dataset.
@@ 5.3 Qualitative Results
Figure 4: PCA visualization of the feature maps from different models. Note that PCA is computed on object features only. The inclusion of geometric constraints during training results in fewer high frequency artifacts in the predicted feature maps for our approach.
In Fig.˜4, we visualize PCA projections of object features produced by our model, Geo-SC, and the unsupervised DINO+SD. Our model produces descriptors that vary smoothly over the object surface while uniquely identifying each point. In comparison, the predictions of Geo-SC are noisier, with sudden discontinuities (e.g., bus) and uniform descriptors on regions that have no keypoints (e.g., the body of the cat and the cow). Meanwhile, the unsupervised features fail to separate repeated parts (e.g., plane engines) and produces noisy features in textureless areas (e.g., tv).
Geo-SC
Ours
Figure 5: Visualization of keypoint matches for randomly selected object points. On each source image (left) we randomly sample points on the object of interest and compute their match on the target (right). Colored lines are used as a way to distinguish the points.
In Fig.˜5 we further qualitatively evaluate our model’s ability to generalize to unseen points. We randomly sample points on the source object and compute their matches on the target. Compared with Geo-SC, our approach exhibits better robustness against matching outside the target object, as well as better spatial awareness of points (e.g., Geo-SC matching points from the bottom of the source bus to the roof of the target), leading to higher matching quality. Further examples, along with comparisons to the unsupervised DINOv2+SD backbone, are provided in Fig.˜A4.
Following wandel2025semalign3d, we visualize our learned canonical shapes in Fig.˜6 by collecting predicted canonical coordinates of multiple objects in order to overlap their partial point clouds over the training data. We observe that the spatial organization of $\mathcal{P}$, i.e., the large bold points, captures the general shape of the category, and the predicted coordinates densely span the object surface. Interestingly, our parametrization of the canonical shape Eq.˜5 forces predicted coordinates to lie within the convex hull of $\mathcal{P}$, which explains the incomplete wheel on the motorbike. We also observe that very few points are mapped towards the end of the train, which we attribute to the varying length of trains across instances and the bias toward frontal viewpoints. Note that contrary to wandel2025semalign3d, these are simply visualizations and are not used for inference, meaning this limitation is unlikely to significantly affect performance.
@@ 6 Limitations
While augmenting the test set of SPair-71k with new keypoints enables us to evaluate existing techniques with their provided models and code, our proposed benchmark SPair-U inherits some drawbacks from SPair-71k. In particular, the test set is small, consisting of only 481 images but with over 8,000 test pairs, and the categories are restricted to only common objects and animals. Furthermore, some categories were already labeled with a high number of keypoints where it is potentially easier to detect the newly added ones by relating them to the existing ones. While our findings related to generalization issues in supervised SC techniques remain valid, a larger, higher-quality held-out set of images and keypoints would be beneficial for more extensive evaluation.
@@
In Eq.˜3, we assume there exist a global rigid transformation between the posed keypoints and their canonical counterpart, which in practice is not the case especially for deformable objects. However, the sole purpose of this step is to optimize $\mathcal{P}$ into a coarse spatial organization of 3D keypoints (e.g., making sure that the left hand keypoint generally sits opposite of the right one) in order to allow the computation of local geometric alignment in Eq.˜6. We show in Fig.˜6 and Section˜5.2 that despite this coarse assumption, our method is able to recover a reasonable 3D structure and performs well on deformable objects.
Finally, our assumption that geometry is a good proxy for semantics breaks down for complex object categories with diverse spatial part configurations. For example, cabinets may have different numbers of doors that open in various directions, leading to inconsistent placement of features like handles. Finally, we do not foresee any negative social impacts of our work.
@@ 7 Conclusion
We addressed the challenge of estimating semantic correspondences across different image instances of the same object category. Although recent supervised methods perform well on keypoints seen during training, we show that they often struggle to generalize to unseen keypoints. To overcome this, we introduced a new approach that incorporates geometric constraints during training by learning a continuous canonical manifold specific to each category. Our method outperforms both supervised and unsupervised baselines, as demonstrated on SPair-U – a new dataset we introduce with additional keypoint annotations for the widely used SPair-71k benchmark.
Acknowledgements. HB was supported by the EPSRC Visual AI grant EP/T028572/1.